INAF queue observing, 2018 September 27-28

Observer: R. Carini, A. Gargiulo
Telescope Operator: D. G. Huerta
Support Astronomer: O. Kuhn on call
Instrument: LUCI

Summary

At the sunset the sky is clear
instruments: MODS2 red grating is not working
LBC-blue has some issues LUCI1 AO doesn't work.

Night duration (12\xB0 el twilight) 2:30- 12:15 UT = 9 h 45m

Observing time: 6 40m (68%)

Weather time loss: h m (0%)

technical time loss 3h 5m (32%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
telluric Cresci B HIP92316 luci1 +luci2 ls,   luci1.0045-0048, luci2.0046-0049  
Cresci B J2054 LUCI1+LUCI1 ls, zjspec 2.00/2.00 luci1.0053-0076, luci2.0054-0077 completed
Cresci B J2054 LUCI1+LUCI2 ima 0.07/0.07 luci1.0077-0083, luci2.0078-0084 completed
telluric Polletta HIP 106817 LUCI2 mods, zjSpec   luci1.20180927.0089-0091 (slit1), luci2.0097 (slit2), luci2.0101-0106 (slit1, slit2)  
Polletta PLCK LUCI1+LUCI2 mos, zjSpec 2.6/4.0 h luci1.20180927.0141-0180, luci2.20180927.0111-0150 luci1 images to be checked
telluric Polletta   LUCI1 , zjSpec   luci1.0192-0194 slit 2
Nisini field3 luci1+luic2 mos, K   luci1.0210-0213, luci2.0167-0170 to be repeated

Details

The ISA reports that a Field Stop Misalignment has involved the Nisini images of field2. Since we acquired the calibration for this field two days ago, they can be shifted with respect to the science. We acquire new calibrations (flat + arc) for H and J filter. We do not acquire the calibration for the K filter, since we need to open the dome to perform the alignment of the field stop in time for the sunset. During the procedure, we find that the telescope was probably aligned during the day, so also the calibration taken today could be shifted with respect the science. For the reduction, please check which are the best.

Calibration Nisini (flat + arc) field 2
  • luci1.20180927.0011-0024, luci2.20180927.0011-0024 H
  • luci1.20180927.0025-0038, luci2.20180927.0025-0038 J

02:18 Cresci, telluric B

  • luci1.20180927.0042-0044, luci2.20180927.0042-0045 acq
  • luci1.20180927.0045-0048, luci2.20180927.0046-0049 sci

02:36 Cresci, target B, spec

  • 02:41 seeing on dimm 0.98"
  • luci1.20180927.0049-0052, luci2.20180927.0042-0053 acq
  • 02:51 seeing on dimm 1.0"
  • 03:04 seeing on dimm 0.9"
  • 03:30 seeing on dimm 1.1"
  • 03:45 seeing on dimm 0.8"
  • 04:01 seeing on dimm 0.8"
  • luci1.20180927.0053-0076, luci2.20180927.0054-0077 sci

04:08 Cresci, target B, ima
  • luci1.20180927.0077-0083, luci2.20180927.0078-0084 sci

04:17 Polletta, tellurica HIP 106817
  • 04:18 seeing 1.1"
  • 04:23 the telescope doesn't take the preset (we are still on Cresci field, images luci1.20180927.0074-0085, luci2.20180927.0085-0086)
  • 04:25 we send the script again
  • luci1.20180927.0086-0088, luci2.20180927.0087-0089 acq
  • 04:39 seeing on dimm 1.0"
  • 04:44 offset to move in the second slit has failed in telluric science
  • luci1.20180927.0089-0091, luci2.20180927.0090-0092 sci for telluric in the "first" slit. David and Olga try to solve the problem of the offset, but in the end we decide that the best solution is to start the acq from scratch
  • 05:03 we resend the script
  • luci1.20180927.0092-0094, luci2.20180927.0093-0095 acq
  • 05:07 we start the integration again and skip to telluric in the "second" slit. It fails again. The problem is with the guide star of LUCI1. Olga suggest to reload the script changing the telescope mode in TRACK.
  • 05:32 we send the script again. It fails again. We decide to use just LUCI 2.
  • 05:40 we send the script just for LUCI2
  • luci2.20180927.0098-00100 acq
  • luci2.20180927.0101-0106 sci

06:03 Polletta, target PLCK
  • luci1.20180927.0097-00100, luci2.20180927.0107-0110 acq
  • 06:18 Aladin stucks. We need to close and re-open the luci execution panel
  • 06:30 luci1 centering has some problems. We try to find the best centering solution by hand.
  • 06:53 we start integration
  • 07:01 seeing 0.7" (not on dimm)
  • luci1.20180927.0141-0160, luci2.20180927.0111-0130 sci
  • 07:10 seeing on dimm 0.9''
  • 07:13 some burst of seeing to 1.1''', but the mean seeing is 0.9'''
  • 07:26 seeing 0.57'' on dimm
  • 07:46 load again the script
  • luci1.0161-0180, luci2.0131-0150 sci
  • 08:11 seeing 0.8''-0.9''
  • 08:38 during the last 2 images the seeing is about 1.0''

08:41 try again the telluric with only LUCI1
  • 08:45 error during the preset
  • 08:46 send again the script HIP 116886, we try to send the offset manually
  • acq luci1.0181-0183
  • sci luci1.0184-0188
  • the offset is too big, the guider loses the star. the offset manually failed
  • we have written a script for this telluric only for the second slit
  • acq luci1.0189-191
  • sci luci1.0192-0194

09:17 switch to Nisini program, telluric
  • TO collimates
  • 09:28 send the script
  • 09:30 error while try to return mask to the cabinet at MOS
  • 10:12 problem fixed by Dave Thompson

10:13 send again the script
  • acq luci1.0195-0197,luci2.0151-0153
  • seeing about 1.0''
  • after the offset to the other slit: error in the grating.
  • we put manually the correct wavelength, but it doesn't work

10:53 the target is too high, we pass to the field3 and come back to the telluric later.
  • acquisition failed. the mask could have some problems.

11:23 try field3
  • acq luci1.0207-0209, luci2.0163-0166
  • during the luci2 acquisition there were problems, the RTDP calculated the wrong offset
  • 12:02 start the scientific integration
  • sci luci1.0210-0213, luci2.0167-0170

12:19 slew to the telluric
  • 12:21 check pointing
  • 12:25 send again the script
  • preset failed, send again
  • acq luci1.0214-, luci2.0170-=
  • problem with the filter wheel only in luci2. we stop and run again the script but the sun shines. The TO have to close the dome.

Calibration
  • Arc+Flat Polletta
  • luci1.0223-236, luci2.0174-187
  • Arc+Flat Antoniucci
  • luci1.0237-252, luci2.0188-203 Hkspec filter
  • luci1.0253-0266, luci2.0204-0217 zjspec filter
  • Arc Nisini field 3: filter K lamp Xe luci1.0267-0268,luci2.0218-0219

-- %USERSIG{RobertaCarini - 2018-09-27}%

Comments

 
Topic revision: r10 - 01 Oct 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback